# Taking quantisation seriously: a farewell to waves

### v2

@@ 4 Evidence supporting Duane’s model
In addition, it has also been shown that van der Waal’s interactions during diffraction contribute to the pattern, in a manner degenerate with the slit width [49] (similar phenomena are used to argue against wave-particle duality in [16]). In these experiments, beams of molecules are diffracted by silicon nitride gratings. The diffraction patterns are found to depend on the dipole polarisability of the incident molecules. As previously mentioned, the effect of the extra interactions produces a modified “effective slit width”. This is very suggestive of Duane’s explanation, which connects slit width to interactions already. Interestingly, these additional interactions also lift the zeroes of the diffraction pattern. This is unsurprising if diffraction momentum changes are being supplied by interactions in general, as the van der Waal interactions can allow deflection angles otherwise forbidden by the intrinsic lattice momenta. On the other hand, it is quite mystifying if this is actually wave interference at work. The wave boundary conditions are not being altered, yet the interference patterns change. More crucially, this case supplies direct evidence that local interactions within diffraction gratings can supply an effect that is degenerate with the boundary conditions of apparent wave interference, precisely what is being claimed by Duane.
It is evident that we can discuss all of our diffraction scenarios in terms of particle interactions. However, we are now in a situation where the necessity of particles having wave properties is contingent on classical descriptions of the diffraction element. Additionally, there are scenarios where this description cannot avoid inclusion of particle interactions. If we consider the fully quantum description more fundamental, this reduces the wave explanation to an emergent consequence of quantisation, with only heuristic value. Our next step will be to introduce further evidence that only a particle-particle interaction description is physically consistent.
5 The angular double slit
@@
$P(p_{y})\propto\mathrm{sinc}^{2}\left[\frac{\sigma p_{y}}{2\hbar}\right]\cos^{2}\left[\frac{p_{y}\Delta y}{2\hbar}\right]\;,$ (6)
where $p_{y}$ is the transverse momentum, $\sigma$ is the slit width, and $\Delta y$ is the spacing. Writing $\alpha=f\pi$, a simple correspondence can be established between the two by introducing the unit-free momentum $\tilde{p}=2\frac{p_{y}\Delta y}{fh}$ so that
$P(\tilde{p})\propto\mathrm{sinc}^{2}\left[\frac{f\pi\tilde{p}}{2}\frac{\sigma}{\Delta y}\right]\cos^{2}\left[\frac{f\pi\tilde{p}}{2}\right]\;.$ (7)
@@
A vital aspect of this experiment is that $\ell$ modes are “free-space modes” [51]. Meaning they do not change $\ell$ while propagating in free space. Thus, a beam’s angular momentum is set by a discrete exchange at the spatial light modulator and not determined at some point of overlap at the detector. This is consistent with the ability to perform the experiment with classical beams. Thus, our $\ell$ spectrum can be simply interpreted as the fraction of beam power propagating in mode $\ell$.
@@
Figure 1: OAM spectrum from angular double slit when $\alpha=\pi$, $\beta=\frac{\pi}{3}$ plotted as a bar graph. The line plot shows a linear double-slit momentum distribution in $\tilde{p}$ when $\frac{\Delta y}{\sigma}=3$ and $f=1$.
@@ 6 The double slit revisited
An intriguing aspect of the OAM double slit in section 5 is that it does not matter (in principle) where we measure the $\ell$ spectrum. Both near-field and far-field diffraction regimes yield the same spectrum. For wave-based explanations this result is extraordinary. The “interference” pattern we see in the far-field is already imprinted in $\ell$ (or by analogy $p_{y}$) within the near-field. It is easy to demonstrate that this is so. Consider a position wavefunction at the double slit position $\psi(\vec{x})$. This can be written in the momentum representation in terms of Fourier coefficients $a_{k}$ (for each momentum $k$). The momentum spectrum is then $|a_{k}|^{2}$. In this same plane wave basis, propagation in free space merely accumulates a phase factor, or $\psi(\vec{x}+\vec{\delta x})=\int a_{k}\mathrm{e}^{i\vec{k}\cdot\vec{x}}\mathrm{e}^{i\vec{k}\cdot\vec{\delta x}}dk$ after propagating $\vec{\delta x}$. Thus, the Fourier coefficients, and hence the $p$ spectrum, are unchanged by free-space propagation. The same argument holds for the angular case when using, for instance, the Laguerre-Gauss basis [51]. The fact that changes occur in the position space pattern, but not the momentum spectrum is then down to each mode accumulating differing phases. We will revisit this from a particle point of view later.
What the above argument demonstrates is that a consistent momentum distribution undergirds interference patterns, rather than requiring the pattern be determined by wave-overlap at the point of detection. To make this complete, we can demonstrate that the quantum formalism itself requires that the momentum spectrum is determined before wave over-lap can even occur.
We start by formulating a simple, position-representation amplitude for particles passing in the $x$-direction through the linear slit apparatus (positioned at $x=0$ and extending symmetrically in the $y$-direction)
$\psi(y)\propto\Pi\left(\frac{y-\Delta y/2}{\sigma}\right)+\Pi\left(\frac{y+\Delta y/2}{\sigma}\right)$ (8)
@@
$\phi(p_{y})\propto\int_{-\infty}^{\infty}\left[\Pi\left(\frac{y-\Delta y/2}{\sigma}\right)+\Pi\left(\frac{y+\Delta y/2}{\sigma}\right)\right]\mathrm{e}^{\left(-i\frac{\vec{x}\cdot\vec{p}}{\hbar}\right)}dy\;.$ (9)
@@
$\phi(p_{y})\propto\mathrm{sinc}\left[\frac{\sigma p_{y}}{2\hbar}\right]\left(2+\mathrm{e}^{i\frac{\Delta yp_{y}}{\hbar}}+\mathrm{e}^{-i\frac{\Delta yp_{y}}{\hbar}}\right)\;,$ (10)
@@
$|\phi(p_{y})|^{2}\propto\mathrm{sinc}^{2}\left[\frac{\sigma p_{y}}{2\hbar}\right]\cos^{2}\left[\frac{p_{y}\Delta y}{2\hbar}\right]\;.$ (11)
@@
This indicates that the $p_{y}$ spectrum, with both single and double-slit factors, is already imposed at $x=0$, before waves originating from each slit get a chance to overlap. Thus, actual spatial wave-overlap and interference is irrelevant in forming the far-field position distribution, which precisely mirrors $P(p_{y})$ above. This conforms nicely with the $\ell$ analogy, as the beams in question do not change $\ell$ while propagating in free space (note the experiment can be done entirely classically). Meaning their $\ell$ must only be determined at the spatial light modulator, not in some interference region at the detection point (unless one wishes to argue the detector itself couples different modes based on its distance to the slits). Moreover, the maxima of $\cos^{2}\left[\frac{p_{y}\Delta y}{2\hbar}\right]$ occur when $p_{y}=\frac{nh}{\Delta y}$, the same quantisation rule observed by Duane, as $\Delta y$ is the symmetry scale of the slits. Although this is a reflection symmetry, rather than a full periodicity, momentum quantisation appears following similar arguments [22]. Additionally, the factor $\mathrm{sinc}^{2}\left[\frac{\sigma p_{y}}{2\hbar}\right]$ has maxima where $p_{y}=\frac{mh}{2\sigma}$ where $m$ is an odd integer. These maxima are spaced out by integer values of $\frac{h}{\sigma}$ and the minima occur $\frac{h}{2\sigma}$ from the maxima. Since $\sigma$ represents a second symmetry scale, the presence of two simultaneous quantisation rules is unsurprising. The “missing orders” occur when minima $\frac{m^{\prime}h}{\sigma}$, $m^{\prime}\in\mathbb{Z}$ match maxima $\frac{nh}{\Delta y}$. If we let $\sigma=f\Delta y$ this matching condition means when $\frac{m^{\prime}}{f}$ is an integer we find a “missing order” that was disallowed by the $\sigma$ scale quantisation.
What then is the particle equivalent of the mode-dependent phase accumulated during propagation? For waves, this phase determines the variation of the position-space image with distance between the slits and screen. The obvious answer is to consider whether the particle’s position in the slits affects its destination. This would be apparent in the near-field, but averaged out in the far. Thus, we could simply make geometric projections for particles originating from different points in the slits, each sampled from the same momentum distribution above. However, we will not reclaim the full detail of a Fresnel pattern. This is because the derivation above does not expose the details of interactions between incident particles and slits [24]. In that, there is no explicit in-slit position dependence. We should expect particles passing closer to the edges have differing interaction probability with each edge and will contribute to the spectrum differently to a particle passing exactly between both edges. Thus, our above result is the correct total spectrum (integrated over the slit width), but lacks the detail that would result in something of similar complexity to a Fresnel integral for the near-field diffraction pattern. The different propagation phase for each wave mode is thus tracing the in-slit position association of each mode. This also provides a very natural explanation of why diffraction images converge to a far-field form, as this is the long-distance geometric projection of trajectories sampled from the momentum spectrum.
@@
It is clear then that the interference-like behaviour in the diffraction of quantum systems can be consistently viewed as originating from the quantisation of particle momenta. This is evidenced directly by the angular double slit, as well as other experiments invoked in section 4. Additionally, the momentum spectrum is determined at the slits and then remains unchanged during propagation (as shown by the far-field pattern reproducing it) in complete harmony with Duane’s hypothesis. This holds even when considered within the mathematics usually attributed to the wave nature of quantum objects. Notably, there is no amplitude interference in the position picture, but it is present in that of momentum. The Fourier transform is viewed here as an assessment of the global periodic structure of the scatterer, rather than a collection of interfering plane waves. The status of momentum can be further reinforced by its conservation for individual recoil events in double slit systems [52]. Such an effect could only occur retrocausally in any interpretation where the interaction occurs via waves. That is, after the particle’s momentum is determined by capture at the screen, but is perfectly natural in Duane’s picture.
7 Coherence and “which way” information
The double slit experiment requires light which exhibits spatial coherence, it is natural to ask how this requirement is interpreted without invoking waves. The coherence between two waves or fields is commonly described as their statistical similarity [53]. This is reinforced for quantum coherence as there is some correspondence between it and entanglement [54, 55]. This has a very natural interpretation in the language of particles: we need a particle distribution whose momentum statistics will not obscure the effect of the double-slit scattering itself. More generally, a multimode photon source has a coherence length because differences in photon momenta become apparent over a long enough distance, hence the dependence of a coherence length on bandwidth and the speed of light. In a diffraction scenario this would produce blurring via incompletely overlapping patterns for each mode.
Another interesting case is if we separate the double slits far enough apart, the cosine squared in Eq. (11) oscillates rapidly between $1$ and 0 with the overall pattern being two single slit diffraction results. This is sometimes discussed as an effect of coherence. However, we see it is because $\frac{h}{\Delta y}\ll p_{y}$, i.e. it is a continuum limit for the double-slit quantisation where the fringe spacing goes to 0 compared to the single slits. The comparison to $p_{y}$ explains why the wavelength is often compared to $\Delta y$ to determine an appropriate slit separation to see double-slit interference, as we can reformulate our condition to $\frac{h}{\Delta y}\ll\frac{h\sin\theta}{\lambda}$ or $\Delta y\gg\lambda$ (as $\sin\theta$ is bounded).
@@ 7 Coherence and “which way” information
A final point of interest is “which way” information. This is a phenomenon where determining which slit a scattered particle traversed can remove the interference pattern [1]. More interestingly, an uncertain determination of the path merely reduces the contrast of the pattern in proportion to the uncertainty [56]. This appears to be strong argument for wave behaviour, as by ‘marking’ the waves emanating from different slits, we are removing the essential coherence needed to produce interference. However, it is very natural in particle language as well. To illustrate, we will consider the case of determining “which path” information by scattering photons off of particles as they emerge from the slits. We choose this case for physical transparency. It is well known that the spatial resolution of such scattering detections is limited by probe wavelength. However, this can be easily proven for a pure particle approach. All we need do is regard a particle emerging form the slits and absorbing a probe photon as being excited. This excitation has an average lifetime $\tau$ that obeys $E\cdot\tau\geq\frac{\hbar}{2}$ where $E$ is the excitation energy. Thus, the position uncertainty of this probe is the distance moved as a result of absorption of the probe photon prior to re-emission, or $\Delta x=\Delta v\tau$. Here $\Delta v$ is the velocity kick introduced by absorption. Now we can write
@@
Where the second equality assumed that we can ignore the initial trajectory of the particle (we are only interested in the excitation and can happily choose the initial rest-frame of the particle). Since $m\Delta v=p_{\gamma}$, or the probe momentum, we have that $\Delta x\geq\frac{\hbar}{p_{\gamma}}$. Now we can see that discriminating between slits will require photons with momenta $\sim\frac{h}{\Delta y}$, very similar to the momentum quantisation responsible for the apparent interference pattern. Moreover, we can imagine that the particle’s $p_{y}$ is randomly changed during the probe scattering, as there are no preferential scattering angles. We can implement this by convolving $P(p_{y})$ with a flat window-function between $p_{y}-\frac{h}{\alpha\Delta y}$ and $p_{y}+\frac{h}{\alpha\Delta y}$ where $\alpha\in\mathbb{R}$. Higher resolution, i.e. $\alpha<1$, results in a broader convolution that reduces the contrast of the bright and dark fringes.

### v4

@@ 5.3 Diffraction of polarisable molecules
Experiments with polarisable molecules demonstrate that van der Waal’s interactions during diffraction contribute to the pattern, in a manner degenerate with the slit width [41] (similar phenomena are used to argue against wave-particle duality in [42]). In these experiments, beams of molecules are diffracted by silicon nitride gratings. The diffraction patterns are found to depend on the dipole polarisability of the incident molecules. As previously mentioned, the effect of the extra interactions produces a modified “effective slit width”. This is very suggestive of Duane’s explanation, which connects slit width to interactions already. It should be noted that further evidence of interactions contributing to the diffraction pattern can be found in [43].
Interestingly, these additional interactions also lift the zeroes of the diffraction pattern. This is unsurprising if diffraction momentum changes are being supplied by interactions in general, as the van der Waal interactions can allow deflection angles otherwise forbidden by the intrinsic lattice momenta. More crucially, this case supplies direct evidence that local interactions within diffraction gratings can supply an effect that is degenerate with the boundary conditions of apparent wave interference, precisely what is being claimed by Duane.
5.4 Coherence
Diffraction experiments like the double slit are often stated to require “spatial coherence” of incoming particles. However, this is severely questionable, as [44] shows that only collimation is needed. Indeed, coherent and incoherent waves are indistinguishable, provided they have negligible dispersion in perpendicular momentum on the scales relevant to the experiment. This already seems to be both powerful evidence for the particle picture presented here and against a wave explanation.
The collimation requirement makes complete sense for a particle picture. Double slit interference requires an incident particle distribution whose momentum statistics will not obscure the effect of the double-slit scattering itself. If one wishes to preserve a wave view, the only explanation can be that coherence is imposed on incoherent incoming waves by their interaction with the slits. This is very difficult to square with the slits behaving as a classical, inert barrier. Clearly, the interaction requirement is very consistent with the notion that wave explanations emerge from ignoring interactions, as is similarly argued in [14].
More generally, a multimode photon source has a coherence length because differences in photon momenta become apparent over a long enough distance, hence the dependence of a coherence length on bandwidth ($\delta p_{\parallel}$ in the direction of motion) and the speed of light. In a diffraction scenario this would produce blurring via incompletely overlapping patterns for each mode, as their deflection angles will all differ due to dependence on $\frac{p_{\perp}}{p_{\parallel}}$.
Another interesting aspect of coherence is that longitudinal coherence lengths do not change as a quantum wave packet spreads [45]. This can lead to situations where waves overlap but do not interfere, an extremely counterintuitive, yet experimentally confirmed [46], result for wave explanations of quantum phenomena. In some scenarios the coherence length and size of the wave packet can even become completely unrelated [47]. This teeters on the brink of total contradiction.
5.5 Momentum conservation
It is commonly asserted that conservation laws in quantum mechanics work only at the statistical level. This being due to the fact that the formalism can only provide such rules for the averages of conserved observables. A particular issue is that of superpositions. If a superposition collapses then there is some non-conservation occurring.
A particle viewpoint, as suggested by Duane, requires non-statistical conservation of both momentum and energy. Importantly, experimental observations of recoiling double-slit analogues [48] require an assumption of momentum conservation on the individual particle level [43], in order for the conventional explanation to succeed. Additionally, single particle orbital angular momentum conservation was observed in [49]. Finally, there exists a theoretical justification to generalise this in [50, 51].
If momentum is indeed individually conserved in double slit scenarios, it is miraculous for wave explanations. It requires that the slit recoil is determined retrocausally. That is, after the particle’s momentum is determined by capture at the screen. However, such conservation is perfectly natural in Duane’s picture.
5.6 The angular double slit
@@
$P(p_{y})\propto\mathrm{sinc}^{2}\left[\frac{\sigma p_{y}}{2\hbar}\right]\cos^{2}\left[\frac{\Delta_{y}p_{y}}{2\hbar}\right]\;,$ (8)
where $p_{y}$ is the transverse momentum, $\sigma$ is the slit width, and $\Delta_{y}$ is the spacing. Writing $\alpha=f\pi$, a simple correspondence can be established between the two by introducing the unit-free momentum $\tilde{p}=2\frac{\Delta_{y}p_{y}}{fh}$ so that
$P(\tilde{p})\propto\mathrm{sinc}^{2}\left[\frac{f\pi\tilde{p}}{2}\frac{\sigma}{\Delta_{y}}\right]\cos^{2}\left[\frac{f\pi\tilde{p}}{2}\right]\;.$ (9)
@@
A vital aspect of this experiment is that $\ell$ modes are “free-space modes” [53]. Meaning a beam does not change $\ell$ while propagating in free space. Thus, a beam’s angular momentum is set by a discrete exchange at the spatial light modulator and not determined at some point of overlap at the detector. This is consistent with the ability to perform the experiment with classical beams. So, our $\ell$ spectrum can be simply interpreted as the fraction of beam power propagating in mode $\ell$.
@@
Figure 1: OAM spectrum from angular double slit when $\alpha=\pi$, $\beta=\frac{\pi}{3}$ plotted as a bar graph. The line plot shows a linear double-slit momentum distribution in $\tilde{p}$ when $\frac{\Delta_{y}}{\sigma}=3$ and $f=1$.
@@ 6 The double slit revisited
An intriguing aspect of the OAM double slit in section 5.6 is that it does not matter (in principle) where we measure the OAM spectrum. Both near-field and far-field diffraction regimes yield the same spectrum. For wave-based explanations this result is extraordinary. The “interference” pattern we see in the far-field is already imprinted in $\ell$ (or by analogy $p_{y}$) within the near-field. It is easy to demonstrate that this is so. Consider a position wavefunction at the double slit position $\psi(\vec{x})$. This can be written in the momentum representation in terms of Fourier coefficients $a_{k}$ (for each momentum $k$). The momentum spectrum is then $|a_{k}|^{2}$. In this same plane wave basis, propagation in free space merely accumulates a phase factor, or $a_{k}e^{i\vec{k}\cdot\vec{x}}\to a_{k}e^{i\vec{k}\cdot\vec{x}}e^{i\vec{k}\cdot\vec{\delta x}}$ after propagating $\vec{\delta x}$ from $\vec{x}$. Thus, the Fourier coefficients, and hence the $p$ spectrum, are unchanged by free-space propagation. The same argument holds for the angular case when using, for instance, the Laguerre-Gauss basis [53]. The fact that changes occur in the position space pattern, but not the momentum spectrum is then down to each mode accumulating differing phases (Gouy phase). We will revisit this from a particle point of view later.
What the above argument demonstrates is that a consistent momentum distribution undergirds interference patterns, rather than requiring particle momenta be determined by wave-overlap at the point of detection. In fact, the Gouy phase also demonstrates each quantum object has a consistent momentum, i.e. it cannot possess a history without a consistent trajectory. Otherwise, it would accumulate a mixed Gouy phase by oscillating in $p$ during propagation. To make this complete, we can demonstrate that the quantum formalism itself requires that the momentum spectrum is determined before wave over-lap can even occur.
6.1 Double slit momentum distributions
We start by formulating a simple, position-representation amplitude for particles passing in the $x$-direction through the linear slit apparatus. This is positioned at $x=0$ and extends symmetrically in the $y$-direction. Thus, we use the normalisable state
$\psi(y)\propto\Pi\left(\frac{y-\Delta_{y}/2}{\sigma}\right)+\Pi\left(\frac{y+\Delta_{y}/2}{\sigma}\right)\;,$ (10)
@@ 6.1 Double slit momentum distributions
$\phi(p_{y})\propto\int_{-\infty}^{\infty}\psi(y)e^{\left(-i\frac{\vec{x}\cdot\vec{p}}{\hbar}\right)}dy\;.$ (11)
@@
$\phi(p_{y})\propto\mathrm{sinc}\left[\frac{\sigma p_{y}}{2\hbar}\right]\left(e^{i\frac{\Delta_{y}p_{y}}{2\hbar}}+e^{-i\frac{\Delta_{y}p_{y}}{2\hbar}}\right)\;,$ (12)
@@
$|\phi(p_{y})|^{2}\propto\mathrm{sinc}^{2}\left[\frac{\sigma p_{y}}{2\hbar}\right]\cos^{2}\left[\frac{\Delta_{y}p_{y}}{2\hbar}\right]\;.$ (13)
This indicates that the $p_{y}$ spectrum, with both single and double-slit factors, is already imposed at $x=0$, before waves originating from each slit get a chance to overlap.
Thus, actual spatial wave-overlap and interference is irrelevant in forming the far-field position distribution, which precisely mirrors $P(p_{y})$ above. This conforms nicely with the $\ell$ analogy, as the beams in question do not change $\ell$ while propagating in free space. Meaning their $\ell$ must only be determined at the spatial light modulator, not in some interference region at the detection point (unless one wishes to argue the detector itself couples different modes based on its distance to the slits).
@@
Moreover, the maxima of $\cos^{2}\left[\frac{\Delta_{y}p_{y}}{2\hbar}\right]$ occur when $p_{y}=\frac{nh}{\Delta_{y}}$, the same quantisation rule observed by Duane, as $\Delta_{y}$ is the symmetry scale of the slits. Although this is a limited translation symmetry, rather than a full periodicity, momentum quantisation appears following similar arguments [22]. Additionally, the factor $\mathrm{sinc}^{2}\left[\frac{\sigma p_{y}}{2\hbar}\right]$ has maxima where $p_{y}=\frac{mh}{2\sigma}$ where $m$ is an odd integer. These maxima are spaced out by integer values of $\frac{h}{\sigma}$ and the minima occur $\frac{h}{2\sigma}$ from the maxima. Since $\sigma$ represents a second symmetry scale for reflection around the slit centre, the presence of two simultaneous quantisation rules is unsurprising. The “missing orders” occur when minima $\frac{m^{\prime}h}{\sigma}$, $m^{\prime}\in\mathbb{Z}$ match maxima $\frac{nh}{\Delta_{y}}$. If we let $\sigma=f\Delta_{y}$ this matching condition means when $\frac{m^{\prime}}{f}$ is an integer we find a “missing order” that was disallowed by the $\sigma$ scale quantisation.
@@
It is clear then that the interference-like behaviour in the diffraction of quantum systems can be consistently viewed as originating from the quantisation of particle momenta. This is evidenced directly by the angular double slit, as well as other experiments invoked in section 5. Additionally, the momentum spectrum is determined at the slits and then remains unchanged during propagation (as shown by the far-field pattern reproducing it) in complete harmony with Duane’s hypothesis. This holds even when considered within the mathematics usually attributed to the wave nature of quantum objects.
6.2 Gouy phase
What then is the particle equivalent of the mode-dependent phase accumulated during propagation? For waves, this phase determines the variation of the position-space image with distance between the slits and screen. The obvious answer is to consider whether the particle’s position in the slits affects its destination. This would be apparent in the near-field, but by definition unnoticeable in the far. Thus, we could simply make geometric projections for particles originating from different points in the slits, each sampled from the same momentum distribution above.
However, we will not reclaim the full detail of a Fresnel pattern. This is because the derivation above does not expose the details of interactions between incident particles and slits [24]. In that, there is no explicit in-slit position dependence. We should expect particles passing closer to the edges have differing interaction probability with each edge and will contribute to the spectrum differently to a particle passing exactly between both edges. Our above result is the correct total spectrum (integrated over the slit width), but lacks the detail that would result in something of similar complexity to a Fresnel integral for the near-field diffraction pattern.
We thus suggest the different propagation phase for each wave mode is tracing their in-slit position associations. This also provides a very natural explanation of why diffraction images converge to a far-field form, as this is the long-distance geometric projection of trajectories sampled from the total momentum spectrum.
7 “Which-way” information
Now we turn to “which-way” information. This is a phenomenon where determining which slit a scattered particle traversed can remove the interference pattern [1]. More interestingly, an uncertain determination of the path merely reduces the contrast of the pattern in proportion to the uncertainty [54]. This appears to be strong argument for wave behaviour, as by ‘marking’ the waves emanating from different slits, we are removing the essential coherence needed to produce interference. However, the incoming wave does not need any actual wave coherence [44], so this explanation fails.
@@ 6.2 Gouy phase
We can address this quite simply in “particle language” instead. To illustrate, we will consider the case of determining “which path” information by scattering photons off of particles as they emerge from the slits. We choose this case for physical transparency. It is well known that the spatial resolution of such scattering detections is limited by probe wavelength. However, this can be easily proven for a pure particle approach. All we need do is regard a particle emerging from the slits and absorbing a probe photon as being excited. This excitation has an average lifetime $\tau$ that obeys $E\cdot\tau\geq\frac{\hbar}{2}$ where $E$ is the excitation energy. Thus, the position uncertainty of this probe is the distance moved as a result of absorption of the probe photon prior to re-emission, or $\Delta x=\Delta v\tau$. Here $\Delta v$ is the velocity kick introduced by absorption. Now we can write
@@
Where the second equality assumed that we can ignore the initial trajectory of the particle (we are only interested in the excitation and can happily choose the initial rest-frame of the particle). Since $m\Delta v=p_{\gamma}$, or the probe momentum, we have that $\Delta x\geq\frac{\hbar}{p_{\gamma}}$. Now we can see that discriminating between slits will require photons with momenta $\sim\frac{h}{\Delta_{y}}$, very similar to the momentum quantisation responsible for the apparent interference pattern. Moreover, we can imagine that the particle’s $p_{y}$ is randomly changed during the probe scattering, as there are no preferential scattering angles. We can implement this by convolving $P(p_{y})$ with a flat window-function between $p_{y}-\frac{h}{\alpha\Delta_{y}}$ and $p_{y}+\frac{h}{\alpha\Delta_{y}}$ where $\alpha\in\mathbb{R}$. Higher resolution, i.e. $\alpha<1$, results in a broader convolution that reduces the contrast of the bright and dark fringes.
7.1 Polarised double slits
What about the more experimentally practical “which-way” case? Where photons pass through a double slit where each slit has one of a pair of orthogonal polarisers before or after it. Here the physics is far less transparent. However, we can see that such a situation disrupts the symmetry of our diffraction set up. Thus, we should expect it to alter the outgoing photon momentum. Specifically, we should see some modification to the double slit quantisation effect but not that of the single slits. In the case where photons can only traverse one slit, as a result of the polarisers, we should lose the double slit pattern entirely. This is what gives rise to the apparent need for wave overlap. However, we can see it is a mathematical consequence of lost symmetry. This does nothing to identify physical processes at work, as we have no detail of the interaction between photon and scatterer. However, in the case of modified, rather than destroyed, symmetry we can make more progress.
Consider a double slit being illuminated by unpolarised light. One slit is preceded by a right-handed ($R$) circular polariser, and other left-handed ($L$). This famously yields “no interference”, or, more accurately, only a single-slit diffraction pattern. However, if we interpose a linear polariser before the screen, a full interference pattern magically returns. In fact, there are two patterns, one for each orthogonal linear polarisation [55] (and a range of intermediate combinations for different axis choices). One such pattern has the same form as the conventional double slit. The other has a central point of destructive interference, rather than a maximum. This is a baffling experiment if one is committed to an explanation in terms of interfering waves. The orthogonal, circular polarisations surely prevent meaningful wave overlap, yet a linear polariser restores an interference pattern. If the linear polariser changes our waves to allow overlap again, why does the propagation distance from the slits (not linear polariser) determine the observed pattern? This has lead to somewhat fanciful interpretations in terms of “which-way” information, which is erased by the linear polariser. Furthermore, we cannot invoke “coherence” to provide a more objective wave explanation [44] without admitting that it is interactions that shape this “interference”, rather than waviness.
This remarkable result is far easier to understand in terms of particle interactions. Our two polarisers have opposite handedness, so we can consider an opposite parity double slit position amplitude
$\psi(y)\propto\Pi\left(\frac{y-\Delta_{y}/2}{\sigma}\right)-\Pi\left(\frac{y+\Delta_{y}/2}{\sigma}\right)\;.$ (15)
It can be easily shown that a Fourier transform yields
$|\phi(p_{y})|^{2}\propto\mathrm{sinc}^{2}\left[\frac{\sigma p_{y}}{2\hbar}\right]\sin^{2}\left[\frac{\Delta_{y}p_{y}}{2\hbar}\right]\;.$ (16)
In other words, an odd-parity amplitude leads to a momentum spectrum with a central 0. This trivially adds with the even version to produce $\mathrm{sinc}^{2}\left[\frac{\sigma p_{y}}{2\hbar}\right]$, a pure single slit pattern. Surprisingly, these odd and even states form a single-slit momentum pattern under either amplitude or probability addition. It must be emphasised that the source of any parity differences must be traced to an interaction between photons and the polarisers.
We can now understand our “magic eraser”. Complementary interference patterns are simultaneously visible without the linear polariser. One of a complementary pair is revealed by the linear polariser, regardless of axis choice. This is a critical signature of an interaction explanation, as these interactions are not turned off by the existence of “which-way” information.
The source of the opposite parity patterns can be found if we decompose our circular polarisations into
[…]
Thus, we must conclude that the $\left|H\right$> and $\left|V\right$> states (or any similar orthogonal pair) carry the two different interference patterns seen with this choice of basis for the final polariser. Moreover, we also know each slit emits both interference patterns. These two states having opposite parity in the $\left|R\right$>, $\left|L\right$> basis completes the link.
Our conclusion is that this is still susceptible to our interaction analysis, despite the failure of a wavey coherence attempt. We have modified the global symmetry via the orthogonal polarisers and thus altered the final momentum distribution. The operation of the polarisers alters the photon states, creating correlations between polarisation and momentum, as implied by sharing the same mathematics as a Stern-Gerlach apparatus, although this seems more subtle for photons. These correlations establish the parity difference that translates to the final momentum distribution. Crucially, we see the symmetry is an aspect of both the incoming photons and the scattering slits, this is not revealed until we impose this parity difference on the photons entering the slits. The magic of quantum mechanics is that we can predict the correct outcomes from symmetry principles, even without a characterisation of the particle interactions. This is not mysterious in a particle view, as the outcome of the interactions is constrained by quantisation, which is imposed by the symmetry.